A car covers 200 m while increasing its velocity from 20 m/s o 30 m/s . Compute he acceleration


1
Expert's answer
2019-09-10T13:27:08-0400

The distance traveled by the car


"d=\frac{v_f^2-v_i^2}{2a}"

So, the acceleration of the car

"a=\frac{v_f^2-v_i^2}{2d}=\frac{(30\:\rm m/s)^2-(20\: m/s)^2}{2\times 200\:\rm m}=1.25\:\rm m/s^2"